OpenStudy (anonymous):

What is the remainder when (2x4 – 3x3 + x2 – 3x – 72) ÷ (x – 2)

OpenStudy (anonymous):

could you maybe clarify your multiplication sign and x

OpenStudy (anonymous):

What is the remainder when (2x^4 – 3x^3 + x^2 – 3^x – 72) ÷ (x – 2)

OpenStudy (anonymous):

i think for 3^x you meant 3x...?

OpenStudy (zehanz):

I think it is this: \(\dfrac{2x^4-3x^3+x^2-3x-72}{x-2}\)

OpenStudy (zehanz):

Do you know how to do a synthetic division?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the options are 6 ,3 -6 or-66

OpenStudy (zehanz):

It goes like this: take the solution of x-2=0, so that is 2. Next to it, write down all the coefficients of 2x4 – 3x3 + x2 – 3x – 72, that is: 2, -3, 1, -3 and -72. Write them down like this and drop the first coefficient below the horizontal line: 2 | 2 -3 1 -3 -72 | ------------------ 2

OpenStudy (zehanz):

Now multiply the first 2 with the number below the line (also 2) and put the result (4) in the next column and add: 2 | 2 -3 1 -3 -72 | 4 ------------------ 2 1

OpenStudy (zehanz):

Multiply the first 2 with 1, put the result in the next column and add: 2 | 2 -3 1 -3 -72 | 4 2 ------------------ + 2 1 3 Keep doing this until the end.

OpenStudy (anonymous):

The remainder is -66 by the way, so i let you look if you get the same with your method Zehanz.

OpenStudy (zehanz):

2 | 2 -3 1 -3 -72 | 4 2 6 6 ------------------ + 2 1 3 3 ... The number that will appear on the dots is the remainder!

OpenStudy (zehanz):

So the remainder is: -66. We now could write that division as: \(\dfrac{2x^4-3x^3+x^2-3x-72}{x-2}=2x^3+x^2+3x+3-\dfrac{66}{x-2}\), because the numbers we found below the horizontal line in the synthetic division are the coefficients of the quotient.

OpenStudy (anonymous):

-66

OpenStudy (zehanz):

Of course, we could just substitute x=2 in 2x4 – 3x3 + x2 – 3x – 72, to get -66, but doing the synth dev gives you something extra!